Owl shift report 07/08/99

Valeri Jejer (jejer@fnal.gov)
Thu, 8 Jul 1999 07:19:56 -0500 (CDT)

Shift Summary for owl shift, 08/Jul/99

On Shift: Julie, Valeri

Hours of beam available at all intensities: 5.75
Hours of beam available at nominal intensity: 5.75
Hours of E832 data (special runs not included): 5.25
Hours of special runs (laser scans, muon runs...): 0.25

Accounting of lost time:

0.25 hr stop/start run, ramp up NM4AN

========================================================================
Runs ended this shift:
Run 13836 - 1.4M events - BEAM_MUON

Tapes brought to FCC this shift: none
(including tapes for online split)

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
Waiting for beam, muon run, E832 datataking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
xxxx xxx 00:00 Continued switching to muon run mode which started
at the end of evening shift.

00:05 Loaded QKPE05 into tps2d4 on ktev3 after onsplit
request.

13836 000 00:20 Initializing muon run. Etot maps failed to load,
loaded them from ktrig window.

01:17 Beam has returned.

01:30 Stop muon run.

01:35 Ramp up NM4AN with positive polarity.

13837 001 01:44 Start BEAM_832 run.

02:00 Red Epicure button sometimes appears in SDAQ
window caused by reading NM1U current offspill.

02:10 CSI 1497 ped. RMS = 3.3.

02:17 SDAQ button in SDAQ window turned red asking
Theo or Ashkan to create a new swics DB file.
The request was e-mailed to the both.

036 02:32 Ktevana errors - missing FFFF word from several
stream 5 crates. Kfrend fixed it.

03:30 Noticed that BOA went up to ~1.5. Called MCR,
they fixed it.

06:00 BOA went up again, call MCR again. They said
that the person who was tuning the beam to
improve the BOA has gone on break. They will
continue to work on it.

Average BOA over run = 1.33, RMS = 0.15.

06:31 Gave Val wake-up call.

07:05 No day shift crew and no estimate at this time.

07:15 Hand in run to the day crew.
========================================================================
Follow up Notes on Previous Shift

Muon run after beam has returned.

========================================================================
Items to Follow up on Next Shift

E832 datataking.